OpenStudy (alex6799):

What is the inverse of the given relation? y = 4x - 8 A. y = 4x - 2 B. C. y = 4x + 2 D.

OpenStudy (alex6799):

\[B: y=\frac{ 1 }{ 4 }x-2\] \[D: y=\frac{ 1 }{ 4 }x+2\]